An angle measures 58° more than the measure of a supplementary angle. What is the measure of each angle?
Answer by checkley79(3341)

LET X & X+58 BE THE 2 ANGLES.
X+X+58=180
2X=180-58
2X=122
X=122/2
X=61 DEGREES FOR THE SMALLER ANGLE.
61+58=119 DEGREES FOR THE LARGER ANGLE.
PROOF:
61+119=180
180=180